Of the following, which is less than 1/4 ?
A. 2/5 B. 4/7 C. 4/11 D. 5/11 E. 6/25



Answer by vleith(2825)

The algebraic way to do this is to put each potential solution over a common denominator with 1/4
2/5 = 8/20 1/4 = 5/20 so 2.5 is > 1/4
4/7 = 16/28 1/4 = 7/28 so 4/7 is > 1/4
and so on
6/25 = 24/100 1/4 = 25/100 so 6/25 is < 1/4
Another way to solve this is to use number sense
2/5 is > 2/6 which is 1/3. which is > 1/4. So 2/5 is > as well
4/7 is > 4/8 which is 1/2. again too big
4/11 is > 4/12 which is 1/3. again too big
5/11 is just less than 5/10 or 1/2. again too big
6/25 is just less than 6/24 oe 1/4. so 6/25 is the only choice that works